OpenStudy (yoongilife):

I am willing to work. I'm here to learn! Topic: Algebra II I will post question once thread becomes active. Thanks.

OpenStudy (yoongilife):

Refer to the Probability Distribution Graph. Question: Which Probabilities are equal to 0.2? Possible Answers: P\[(1 \le X \le 3)\] (NO) Would we use the formula for Area of Triangle to calculate this one? A = ½ bh A = ½ (3)(0.3?) A= ½ (0.9) A=0.45 P\[(X \ge 4) \] (SELECTION) A = ½ bh A = ½ (1)(0.4) A= 0.2 P\[(X \le 2)\] (SELECTION) A= ½ bh A= ½ (2)(0.2) A=0.2 P\[(2 \le X \le 4)\] I’m a little unsure on how to go about solving this one. It’s not really acceptable to use the formula for finding the area of a triangle, but a trapezoid formula wouldn’t work either, right? I apologize, but I’m a newbie in the universe of Algebra II.